# Scaling-aware rating of count forecasts

### v4

@@ 3.2 Necessity for imperfect reference forecasts
The setting so far was artificial: The baseline model is ideally Poissonian, calibrated, and unsurprisingly matches the theoretical expectation. We merely confirmed that the Expectation Maximization algorithm of A works. Figure 3 explains why model performance across departments is only ostensibly different. Characterizing the ideal case is, however, not sufficient for judging forecasts in practice: One needs to quantify how “close” the model performance is to ideal, and rate this distance to conclude whether the departure from the ideal case is acceptable or problematic. For this purpose, we need to define imperfect benchmarks that set the standard for “excellent”, “good”, “OK”, “fair”, ”insufficient” and “unacceptable” performance, both regarding the level of bias and the level of noise.
@@
Perfect performance is uniquely defined, there is one way to be “right”: Given a Poisson-noise-limited situation, the best possible value of any metric is the value achieved for outcomes drawn from the Poisson distribution. For the bias (the quotient mean prediction / mean observation), the ideal value is one. For other metrics, analytic formulae are often available (see Wei and Held (2014) for WMAPE/MAE and (N)MRPS); in the worst case, one can recur to numerical computation by Monte Carlo sampling.
@@
There are, however, many ways to be wrong: A forecast can be biased, it can be affected by noise, suffer from both, or by more complicated artefacts (e.g., finite-stock effects, zero-inflation…). No general “worst benchmark model” can be found for unbounded metrics that we are dealing with: Given a forecast, one can construct arbitrarily incompatible (“bad”) observations; given a set of observations, one can construct arbitrarily terrible forecasts. A universal definition of “excellent” to “unacceptable” models that would fit all industries is unlikely to be ever possible (Kolassa, 2008). We therefore expose the industry-specificity by few parameters within a simple model for imperfect reference forecasts. As a “bad” reference forecast, we use a collective global forecast which does not individualize at all, but produces the same value for all products, locations and days.
3.3 Scaling of imperfect reference forecasts
Given predictions of $r_{j}\approx 10$ ($R=1$), we ideally expect the histogram of observations to match the Poisson distribution with rate 10. To simplify judging the deviation from that ideal, we use negative-binomial distributions of different variances to set reference ratings, with examples in Figure 4. That is, we assume that the forecaster produces a Poisson-forecast, but that the observations are generated by a negative-binomial process, such that they observe an over-dispersed distribution of outcomes. We assign the variances at prediction $r_{j}=10$ and the overall bias factors (mean prediction / mean observation) a qualitative score:
@@ 3.3 Scaling of imperfect reference forecasts
The ratings for bias and for variance are independent, that is, a forecast can enjoy Poisson-like variance, but be quite biased, or vice versa. The 12 parameters fully define our rating, they were chosen such that a collective global forecast for the M5 competition data is rated “Insufficient” by most metrics.
@@
Figure 4: Over-dispersed distributions that define the reference forecasts “perfect” (Poissonian, blue), “excellent” (light green), “good” (dark green), etc., for a predicted rate $\mu=10$. When the forecaster predicts a Poisson-distributed target, we rate their performance with respect to a metric by referring to the metric values achieved by the respective negative binomial distributions shown in the figure.
@@
For a given metric $M$, we need to compute reference values $M_{\text{quality}}^{~{}(R)}$ for each quality (“Excellent”, “Good” etc.), and for each bucket of predictions $\vec{r}^{~{}(R)}$. The achieved value $M_{\text{actual}}^{~{}(R)}=M(\vec{r}^{~{}(R)},\vec{s}^{~{}(R)})$ can then be set into that context to rate each individual bucket. When moving to smaller or larger predictions, the question arises how the variance scales: Having established that for rate 10, a variance of around 30 is judged as “good”, which variance is “good” when the prediction is 1, 100, or 10’000? This problem is illustrated in Figure 5, where the distribution of observations for a “good” forecast of 10 is shown on the upper left panel in red and compared to the blue Poisson distribution.
Super-Poissonian variance can scale with the rate in different ways. We parametrize the variance as function of the rate $\mu$ as follows:
$\text{variance}=\mu+f\mu^{\gamma},$ (11)
@@
where $f$ quantifies the strength of the noise. Setting $\gamma=2$ corresponds to quadratic over-dispersion (a Negative Binomial-2-process), and $\gamma=1$ to linear over-dispersion (a Negative Binomial-1-process) (Hilbe, 2012; Cameron and Trivedi, 2005). In general, the variance $\sigma_{\mu}^{2}$ of the underlying distribution of the rate $\mu$, here assumed to be a $\Gamma$-distribution, is inherited by the variance of the resulting count process via $\text{variance}=\mu+\sigma_{\mu}^{2}$ (Karlis and Xekalaki, 2005). The upper right panel of Figure 5 shows the NMRPS as a function of the predicted rate, for different exponents $\gamma=1,1.5,2$. Quadratic $\gamma=2$ is unrealistic as a candidate for “equal quality”: It is too benevolent for large rates and becomes too strict for small rates. $\gamma=1$ exhibits the opposite behavior: It judges large rates strictly and small rates too benevolently. The intermediate value $\gamma=1.5$ provides a reasonable compromise, corroborated by this argument: When summing negative binomial-distributed random numbers, $\gamma=1$ corresponds to uncorrelated noise, $\gamma=2$ to perfectly correlated noise, and $\gamma=1.5$ to partial correlation, which is realistic in the retail setting: The number of visitors of a supermarket has some uncertainty that leads to correlated, product-independent noise on the demand, but also each individual product suffers from uncorrelated, product-dependent noise.
@@
Figure 5: Upper left: Poisson (blue) and negative binomial (red, $\text{variance}=30$) distributions for rate $\mu=10$. Upper right: expected NMRPS as a function of rate, when the variance equals 30 at prediction 10 (corresponding to “good” as defined in Figure 4), for over-dispersion exponents $\gamma=1$ (green), $\gamma=1.5$ (orange), $\gamma=2$ (black). The blue line delimits the ideally achievable (Poissonian) baseline value under over-dispersion 0. Lower left: rate-10-distribution scaled down to rate 1, using linear (green), quadratic (black) and exponent-1.5 (orange) dispersion. Lower right: rate-10-distribution scaled up to rate 100, using linear, quadratic and exponent-1.5-dispersion.
@@
The resulting rate-dependent reference values for the exponent $\gamma=1.5$ are shown in Figure 6. For a given prediction bucket, one reads off what the achieved metric means in terms of model quality (“excellent”, “good”, etc.) by locating the bucket in the respective plots. By assigning “perfect” the numerical score 100% and “unacceptable” the score 0%, and interpolating between reference lines, each bucket gets assigned a numerical score $S^{(R)}$. We will see below that empirical real-life forecasts match the behavior described by $\gamma=1.5$ very well.
@@
A given metric value, e.g., WMAPE=0.6, can be considered “perfect” for a predicted rate of 2, or “unacceptable” for a predicted rate of 100. By applying the method to different sets of data from different industries, we have experienced that “large” buckets (containing many, more than 100, predictions and observations) that are found outside the “insufficient” region are indeed rare and point the user to severe model or data quality issues. The significance of “small” buckets (containing less than about 100 predictions and observations) is, however, questionable: Each metric inherits the randomness of the process, and will therefore be distributed around its expectation value. An unexpected metric value that a small bucket assumes is not necessarily significant, i.e. it can also be due to chance.
@@
Figure 6: Reference values of metrics as a function of predicted rate, for the different scaling-aware quality ratings. We use the variance values defined for $\mu=10$ for the different qualities in Figure 4, and apply the scaling rule Eq. (11) with the exponent $\gamma=1.5$ to infer the variance at rates $\mu\neq 10$. This figure answers the following question: “What value of $M$ is excellent/good/OK… for a prediction $\mu$?”
@@ 3.4 Summarizing bucket-wise to overall ratings
It is desirable to summarize the information contained in a set of buckets by a single number. There are two possibilities to perform this summary: One can aggregate the ratings per bucket via taking the mean of the bucket-wise scores $S^{(R)}$ (weighted by the number of observations per bucket), which yields an overall score $S_{\text{overall}}$ between 0% (all buckets are “unacceptable”) and 100% (all buckets are “perfect”). Alternatively, one can judge the overall aggregated metric, e.g. the overall achieved WMAPE, by setting it into the context of the metric that would have been achieved if the model quality were “excellent” to “good”. In other words, one can compute the expectation value of the metric under reference models of imperfect quality, answering the important question “what values of the metric could I possibly achieve?”, which we had put forward above in Section 1.1.
@@ 3.5 Ostensibly better-than-Poisson measurements
    Finite capacity: If stocks are finite, not all demand peaks are reflected by corresponding sales values, but censored. Fewer events with $s_{j}\gg r_{j}$ will then be observed than were predicted, and, consequently, metrics such as MAE, WMAPE will achieve lower values.
@@
    Systematic over- or under-prediction. When a bucket is systematically over- or under-predicted ($r_{\text{total}}^{~{}(R)}>s_{\text{total}}^{~{}(R)}$ or $r_{\text{total}}^{~{}(R)}<s_{\text{total}}^{~{}(R)}$) the resulting position in the rate-KPI-diagram will be shifted to the left or right, possibly ending at a “better than Poisson” position. A good noise rating for a bucket should only be taken as a sign that the forecast is good if the bucket is also unbiased.
@@
    Overfitting. When rating a model in-sample, a sub-Poissonian performance can witness overfitting. E.g., trivially setting $r_{j}=s_{j}$ gives an implausible WMAPE of 0.
@@
4 Application to M5-competition-models
As a first application, we apply our scaling-aware rating to several models for the validation period of the M5 competition (Makridakis et al., 2022a, b; kaggle, 2022a).
@@ 4 Application to M5-competition-models
    Baseline. The baseline model described in Section 2.6 (details in A) is calibrated and maximally (Poisson) sharp.
@@
To rate the models, we follow the procedure described above in Section 3: We bucket predictions by their value, and plot the bias and the NMRPS for each bucket in Figure 7. We clip small predicted rates at 0.01 (which gives predictions with $r_{j}<0.01$ an advantage because they are under-predicting, on average), and we clip the bias factor at 10 and 1/10 (giving the otherwise even more heavily biased prediction buckets an advantage) to fit into the visual representation. The plots allow us to judge the models and discuss their strengths and weaknesses.
4.1 Model comparison: Rate-stratified representation
@@
The stratified noise plot in the lower panel of Figure 7 clearly separates the five models: The baseline fits the Poisson-ideal perfect line (reproducing Figure 2). For every prediction bucket, LightGBM is more noisy, rated in the “good” range. The Simple-28-day model is, again, slightly worse, populating the “OK” area. The scattered circles follow the parametrized corridors with $\gamma=1.5$ remarkably well – any departure of these corridors is, thus, a forecasting abnormality that should be investigated. The naïve-1-day-ahead model follows with some distance to the others on the boundary to “insufficient”. The largest two prediction buckets of that model leave the bulk – the regression to the mean is especially pronounced for the largest predictions, which correspond to the largest observations on the day before. Not surprisingly, the collective global model exhibits a large degree of noise, close to “unacceptable”.
@@
We summarize Figure 7 by the following scaling-aware ratings (computed following Subsection 3.4):
@@
With Figure 7 and the resulting ratings at hand, one gains the confidence that the LightGBM-model is consistently better than Simple-28-days, across all prediction buckets and in a stable fashion. The rating of bias per bucket is more stringent than the overall bias alone: The Naïve-1-day-ahead method is, by construction, overall unbiased (it uses the sales of the last day before and of the first 27 days of the validation period), but only the bucket with $R\approx 1$ is unbiased, resulting in a “Fair” overall bias performance. The “insufficient” rating of the global collective model reflects its lack of any individualisation among SKUs and days.
4.2 Dataset comparison
@@
Scaling-aware model rating adds more details to model comparison, but often reproduces the verdict of scaling-unaware metrics. For dataset comparison, the added value of scaling-aware rating is more evident: Here, the conclusion $M_{A}<M_{B}\Rightarrow$ “A is better than B” must not be drawn, because it is almost always affected by precision scaling. Scaling-aware rating then helps identify those data-subsets that require special attention, avoiding the naïve scaling trap.
@@
Figure 8: Prediction-bucketed representation of bias (upper panel) and noise (NMRPS, lower panel), for three exemplary departments (LightGBM model). Circle sizes reflect the square root of the total number of sales contained in the respective buckets.
@@
The bucket scatter plot is shown for LightGBM for three selected departments in Figure 8. The following table summarizes the achieved overall ratings, which complements the metrics displayed in Figure 9:
Department $S_{\text{overall}}$(NMRPS) Quality (NMRPS) $S_{\text{overall}}$(Bias) Quality (Bias) Bias factor
Foods 3 64.9% Good 91.3 % Excellent 1.0
Household 1 75.6% Excellent 83.7 % Excellent 1.0249
Hobbies 2 63.5% Good 46.3% OK 0.8553
@@
The comparison of unscaled metrics and scaling-aware ratings per department allows us to appreciate the added insight from the scaling-aware approach: From the WMAPE, MAE, (N)MRPS alone (displayed in Fig. 9), one cannot make any statement about which department is performing better or worse – context is required. Using the rating, and from the lower panel of Figure 8, we see that “Household 1” has less noise than “Foods 3”, and both these departments perform better than “Hobbies 2”. “Foods 3” enjoys a lack of overall bias (quotient 1.0) and an excellent bias rating (91.3%), reflecting the lack of substantial bias in every bucket. The largest predictions in “Household 1” are clearly deteriorated in both the bias and the noise plots. In “Foods 3”, only the small predictions ($r<0.3$) stand out as being under-forecasting, a general minor mis-calibration of the LightGBM model (see upper panel of Figure 7).
@@
The systematic negative bias in “Hobbies 2” is reflected by the mediocre bias score (46.3%), while the rating on NMRPS is benevolent (63.5%) – a systematic negative bias shifts the reference values for noise to larger values. For the overall KPIs, the reference values of normalized metrics (WMAPE and NMRPS) are rather benevolent (such that the achieved KPIs lie in the “Fair” and “OK” range, respectively), while absolute metrics (MAE and MRPS) are more strict (the KPI ends up in the “unacceptable” and “insufficient” range, respectively). When bias is negligible, the ratings of the different departments speak a clear message: “Household 1” performs better than “Foods 3”, while the scaling-unaware MAE, MRPS, WMAPE, NMRPS only provide inconclusive, mixed messages. The contextualization achieved by adding the expected metric values for the different qualities reproduces this message in Figure 9: “Foods 3” finds itself in the “good” range, “Household 1” reaches the edge to “excellent”. Clear operational conclusions follow: One should focus on “Hobbies 2”, then improve “Foods 3” and finally tackle the outliers in “Household 1”.
Figure 9: Overall achieved KPIs for the three departments, and reference KPI values for “excellent”, “good” etc. Due to different selling velocities in the three departments, the reference values differ. An interpretation of the scaling-unaware metrics is misleading, the context added by the reference values is essential. The WMAPE of 57% in “Foods 3” can be considered worse than the WMAPE of 66% in “Household 1”, which is aligned with the visual impression from the scatter plot in the lower panel of Figure 8 as well as with the achieved $S_{\text{overall}}(\text{NMRPS})$ (Subsection 4.2).
5 Application to Demand Edge for Retail predictions at Sainsbury’s

### v5

@@ 3.2 Imperfect reference forecasts
The setting so far was artificial: The baseline model is ideally Poissonian and unsurprisingly matches the theoretical expectation. We merely confirmed that the Expectation Maximization algorithm works. Figure 2 explains why model performance across departments is only ostensibly different. Characterizing the ideal case is, however, not sufficient for judging forecasts in practice: One needs to quantify and rate how close the model performance is to ideal. For this purpose, we define below use-case dependent benchmark forecasts that set the standard for “Excellent”, “Good”, “OK”, “Fair”, ”Insufficient” and “Unacceptable” performance, both regarding the level of bias and the level of noise.
@@
Perfect performance is uniquely defined, there is one way to be “right”: Given a Poisson-noise-limited situation, the best possible value of any metric is the value achieved for outcomes drawn from the Poisson distribution. For the bias (the quotient mean prediction / mean observation), the ideal value is one. For other metrics, analytic formulae are often available (see Wei and Held (2014) for WMAPE/MAE and (N)MRPS); otherwise, one can recur to Monte Carlo sampling.
@@
There are, however, many ways to be wrong: A forecast can be biased, it can be affected by noise, suffer from both, or by more complicated artefacts. No general “worst benchmark model” can be defined for unbounded metrics: Given a forecast, one can construct arbitrarily incompatible (“bad”) observations; given a set of observations, one can construct arbitrarily terrible forecasts.
While there is no unambiguous way to characterize imperfect forecasts, the following parametrization provides meaningful references: The forecaster continues to produce a Poisson-forecast $\mu=r_{j}$, but the observations $s_{j}$ are drawn from negative-binomial distributions fulfilling
$\text{variance}=\mu+f\mu^{\gamma},$ (10)
where the over-dispersion factor $f\geq 0$ reflects the quality of the forecast ($f=0$ is the Poisson baseline). As we argue below in Section 3.3, it is plausible to set $\gamma=1.5$. In other words, the forecaster underestimates their forecast’s uncertainty; they observe an unexpectedly over-dispersed distribution of outcomes. We now assign different scores to different degrees of over-dispersion $f$ in Figure 3. The factors $f$, the variances at prediction $r_{j}=10$ and the maximal overall bias factors (mean prediction / mean observation) that are rated with a certain verdict are summarized in Table 2.
@@
Table 2: Over-dispersion, variance and bias values that are considered “Perfect”, “Excellent”, etc., in a rating scheme that is adapted to the M5 competition. The table is read as follows: For a prediction of $r_{j}=10$, an observed variance between 26 and below 37 is considered “Good”; a variance above 136 is rated “Unacceptable”; an overall bias (mean prediction / mean observation) between 1.2 and below 2 is judged “Fair”.
@@
The 12 parameters (6 over-dispersion and 6 bias factors) fully define our rating, they were chosen such that a naïve collective global forecast for the M5 competition data is rated “Insufficient” by most metrics. We recommend create a rating by defining the “Insufficient” rating in a way that a naïve base model would be within this rating. The ratings for bias and for variance are independent, that is, a forecast can enjoy Poisson-like variance, but be quite biased, or vice versa.
@@
Figure 3: Over-dispersed distributions that define the reference forecasts “perfect” (Poissonian, blue), “excellent” (light green), “good” (dark green), etc., for a predicted rate $\mu=10$. When the forecaster predicts a Poisson-distributed target, we rate their performance with respect to a metric by reference to the metric values achieved by these “excellent”, “good”, etc. negative binomial distributions.
3.3 Scaling of reference metric values
@@
For a given metric $M$, we need to compute reference values $M_{\text{quality}}^{~(R)}$ for each quality (“Excellent”, “Good” etc.), and for each bucket of predictions $\vec{r}^{~(R)}$. The achieved value $M_{\text{actual}}^{~(R)}=M(\vec{r}^{~(R)},\vec{s}^{~(R)})$ can then be set into that context to rate each individual bucket. We have done this exercise for one value of $R$ in Table 2.
@@
When moving to lower or higher predictions, the question arises how the variance should scale when the quality remains the same: Having established that for rate of 10, a variance of around 30 is judged as “good”, which variance is “good” when the prediction is 1, 100, or 10’000? The variance at the other rates is determined by the exponent $\gamma$, illustrated in Figure 4. Setting $\gamma=2$ corresponds to quadratic over-dispersion (a Negative Binomial-2-process), and $\gamma=1$ to linear over-dispersion (a Negative Binomial-1-process) (Hilbe, 2012; Cameron and Trivedi, 2005). The upper right panel of Figure 4 shows the NMRPS as a function of the predicted rate, for different exponents $\gamma=1,1.5,2$. Quadratic $\gamma=2$ (black line) is unrealistic: It is too benevolent for large rates and too strict for small rates. $\gamma=1$ (green line) exhibits the opposite behavior: It judges large rates strictly and small rates too benevolently. The intermediate value $\gamma=1.5$ (orange) provides a reasonable compromise, corroborated by this argument: When summing negative binomial-distributed random numbers, $\gamma=1$ corresponds to uncorrelated noise, $\gamma=2$ to perfectly correlated noise, and $\gamma=1.5$ to partial correlation. This is realistic in the retail setting: The number of visitors of a supermarket has some uncertainty that leads to correlated, product-independent noise on the demand, but also each individual product suffers from uncorrelated, product-dependent noise.
@@
Figure 4: Upper left: Poisson (blue) and negative binomial (red, $\text{variance}=30$) distributions for rate $\mu=10$. Upper right: expected NMRPS as a function of rate, when the variance equals 30 at prediction 10 (corresponding to “good” as defined in Figure 3), for over-dispersion exponents $\gamma=1$ (green), $\gamma=1.5$ (orange), $\gamma=2$ (black). The blue line delimits the ideally achievable (Poissonian) baseline value under over-dispersion 0. Lower left: rate-10-distribution scaled down to rate 1, using linear (green), quadratic (black) and exponent-1.5 (orange) dispersion. Lower right: rate-10-distribution scaled up to rate 100, using linear (green), quadratic (black) and exponent-1.5 (orange) dispersion.
@@
The resulting rate-dependent reference values for the exponent $\gamma=1.5$ are shown in Figure 5. For a given prediction bucket, one reads off what the achieved metric means in terms of model quality (“Excellent”, “Good”, etc.) by locating the bucket in the respective plots. By assigning “perfect” the numerical score 100% and “unacceptable” the score 0%, and interpolating between reference lines, each bucket $R$ gets assigned a score $S^{(R)}$. We will see below that empirical real-life forecasts match the behavior described by $\gamma=1.5$ well.
@@
A given metric value, e.g., NMRPS=0.5, can be considered “perfect” for a predicted rate of 2, or “unacceptable” for a predicted rate of 100. By applying the method to different sets of data from different industries, we have experienced that “large” buckets (containing many, more than 100, predictions and observations) that are found outside the “insufficient” region are indeed rare and point the user to severe model or data quality issues. An unexpected metric value that a small bucket assumes, on the other hand, is not necessarily significant.
@@
Figure 5: Reference values of NMRPS as a function of predicted rate, for the different scaling-aware quality ratings. We use the variance values defined for $\mu=10$ for the different qualities in Figure 3, and apply the scaling rule Eq. (10) with the exponent $\gamma=1.5$ to infer the variance at rates $\mu\neq 10$. This figure answers the following question: “What value of $M$ is excellent/good/OK… for a prediction $\mu$?”
@@ 3.4 Summarizing bucket-wise to overall ratings
It is desirable to summarize the information contained in a set of buckets by a single number. There are two possibilities to perform this summary, discussed in the Supplemental Material: One can aggregate the ratings per bucket via taking the mean of the bucket-wise scores $S^{(R)}$ (weighted by the number of observations per bucket), which yields an overall score $S_{\text{overall}}$ between 0% (all buckets are “unacceptable”) and 100% (all buckets are “perfect”). Alternatively, one can judge the overall aggregated metric, e.g., the overall achieved WMAPE, by setting it into the context of the metric that would have been achieved if the model quality were “excellent” to “good”. In other words, one can compute the expectation value of the metric under reference models of imperfect quality, answering the important question “what values of the metric could I possibly achieve?”, which we had put forward above in Section 1.1.
@@ 3.5 Ostensibly better-than-Poisson measurements
    Finite capacity: If stocks are finite, not all demand peaks are reflected by corresponding sales values, but censored. Fewer events with $s_{j}\gg r_{j}$ will then be observed than were predicted, and, consequently, error metrics achieve lower values.
@@
    Systematic over- or under-prediction: When a bucket is systematically over- or under-predicted ($r_{\text{total}}^{~(R)}>s_{\text{total}}^{~(R)}$ or $r_{\text{total}}^{~(R)}<s_{\text{total}}^{~(R)}$) the resulting position in the rate vs. metric plot will be shifted to the left or right, possibly ending at a “better than Poisson” position. A good noise rating for a bucket should only be taken as a sign that the forecast is good if the bucket is also unbiased.
@@
    Overfitting: When rating a model in-sample, a sub-Poissonian performance can witness overfitting. E.g., trivially setting $r_{j}=s_{j}$ gives an implausible WMAPE of 0.
@@
4 Applications
4.1 M5-competition models
We apply our scaling-aware rating to several models for the validation period of the M5 competition (Makridakis et al., 2022a, b; kaggle, 2022a).
@@ 4.1 M5-competition models
    Baseline. The “perfect” baseline model described in Section 2.5 (details in the Supplemental Material) is calibrated and maximally (Poisson) sharp.
@@
To rate the models, we follow the procedure described above in Section 3: We bucket predictions by their value, and plot the bias and the NMRPS for each bucket in Figure 6. We clip small predicted rates at 0.01 and we clip the bias factor at 10 and 1/10 to fit into the visual representation.
4.1.1 Model comparison: Rate-stratified representation
@@
The stratified noise plot in the lower panel of Figure 6 clearly separates the five models: The baseline fits the Poisson-ideal perfect line (reproducing Figure 1). For every prediction bucket, LightGBM is more noisy, rated in the “good” range. The Simple-28-day model is, again, slightly worse, populating the “OK” area. The scattered circles follow the parametrized corridors with $\gamma=1.5$ remarkably well. The naïve-1-day-ahead model follows with some distance to the others on the boundary to “insufficient”. The largest two prediction buckets of that model leave the bulk – the regression to the mean is especially pronounced for the largest predictions, which correspond to the largest observations on the day before. Not surprisingly, the collective global model exhibits a large degree of noise, close to “unacceptable”.
@@
Table 3 summarizes Figure 6 via the scaling-aware ratings, computed following Subsection 3.4 (see also the Supplemental Material). With Figure 6 and the resulting ratings at hand, one gains the confidence that the LightGBM-model is consistently better than Simple-28-days, across all prediction buckets and in a stable fashion. The rating of bias per bucket is more stringent than the overall bias alone: The Naïve-1-day-ahead method is, by construction, overall unbiased (it uses the sales of the last day before and of the first 27 days of the validation period), but only the bucket with $R\approx 1$ is unbiased, resulting in a “Fair” overall bias performance. The “insufficient” rating of the global collective model reflects its lack of any individualisation among SKUs and days.
@@
Table 3: Scaling-aware ratings for the different models employed on the M5-competition dataset. The ratings summarize the predicted-bucketed representation of Figure 6.
4.1.2 Dataset comparison: Department performance
@@
Scaling-aware model rating adds more details to model comparison, but often reproduces the verdict of scaling-unaware metrics. For dataset comparison, the added value of scaling-aware rating is more evident: Here, the conclusion $M_{A}<M_{B}\Rightarrow$ “A is better than B” must not be drawn, because it is almost always affected by precision scaling.
@@
Figure 7: Prediction-bucketed representation of noise (NMRPS), for three exemplary departments (LightGBM model). Circle sizes reflect the square root of the total number of sales contained in the respective buckets.
@@
The bucket scatter plot is shown for LightGBM for three selected departments in Figure 7. Table 4 summarizes the achieved NMRPS and overall ratings.
Department NMRPS $S_{\text{overall}}$(NMRPS) Quality (NMRPS) $S_{\text{overall}}$(Bias) Quality (Bias) Bias factor
Foods 3 42% 64.9% Good 91.3 % Excellent 1.0
Household 1 47% 75.6% Excellent 83.7 % Excellent 1.0249
Hobbies 2 80% 63.5% Good 46.3% OK 0.8553
Table 4: Scaling-aware ratings on departments level, for the LightGMB model applied to the M5 competition dataset. The ratings summarize the prediction-bucketed representation of Figure 7.
@@
The comparison of unscaled metrics and scaling-aware ratings per department adds substantial insight: From the NMRPS alone, one cannot make any statement about which department is performing better or worse. Using the rating $S_{\text{overall}}(\text{NMRPS})$, and from the lower panel of Figure 7, we see that “Household 1” has less noise than “Foods 3”, and both these departments perform better than “Hobbies 2”. “Foods 3” enjoys a lack of overall bias (quotient 1.0) and an excellent bias rating (91.3%), reflecting the lack of substantial bias in every bucket. The largest predictions in “Household 1” are clearly deteriorated in both the bias and the noise plots. In “Foods 3”, only the small predictions ($r<0.3$) stand out as being under-forecasting, a general minor mis-calibration of the LightGBM model (see upper panel of Figure 6).
@@
The systematic negative bias in “Hobbies 2” is reflected by the mediocre bias score (46.3%), while the rating on NMRPS is benevolent (63.5%) – a systematic negative bias shifts the reference values for noise to larger values. For the overall KPIs, the reference values of normalized metrics (WMAPE and NMRPS) are rather benevolent, while absolute metrics (MAE and MRPS) are more strict (the KPI ends up in the “unacceptable” and “insufficient” range, respectively). When bias is negligible, the ratings of the different departments in Table 4 speak a clear message: “Household 1” performs better than “Foods 3”, while the scaling-unaware MAE, MRPS, WMAPE, NMRPS only provide inconclusive, mixed messages.
4.2 Demand Edge for Retail predictions at Sainsbury’s